### v1

# On polynomial automorphisms of Nagata type

Abstract.
We define a family of polynomial ring homomorphisms generalising the Nagata automorphism, and relate the condition of non-vanishing of their Jacobian determinant with a certain linear partial differential equation. We also give a condition on that family to get wild automorphisms.
Key words and phrases:
polynomial automorphisms, Nagata automorphism, wild automorphisms, linear PDE
2010 Mathematics Subject Classification:
14R10, 13B25, 13F20, 35F05
1. Introduction
In what follows $F$ will be a field of characteristic 00 unless otherwise stated. The polynomials:
$f(x,y,z)$ $=x-2y(zx+y^{2})-z(zx+y^{2})^{2},$
$g(x,y,z)$ $=y+z(zx+y^{2}),$
$h(x,y,z)$ $=z.$
define an automorphism on $F[x,y,z]$ called Nagata automorphism [1]. The determinant of the Jacobian matrix of this automorphism es $1$. In the following paragraph we recall the notion of tame automorphisms of polynomials. Let $1\leq i\leq n$, the $i$th elementary automorphism $\Phi^{(i)}\colon F^{n}\to F^{n}$ is the automorphism defined as
$\Phi^{(i)}(x_{1},\dots,x_{n})=\Big{(}x_{1},\dots,x_{i-1},x_{i}+f(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{n}),x_{i+1},\dots,x_{n}\Big{)}$
and its inverse is the automorphism
${\Phi^{(i)}}^{-1}(x_{1},\dots,x_{n})=\Big{(}x_{1},\dots,x_{i-1},x_{i}-f(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{n}),x_{i+1},\dots,x_{n}\Big{)}\,.$
One may notice that the Jacobian matriz of $\Phi^{(i)}$ has determinant $1$. An automorphism is tame if it is a finite composition of elementary and linear automorphisms. An automorphism that is not tame is called wild. In 1972, Nagata conjectured that the Nagata automorphism is wild. In [2, Corollary 9], the authors proved this conjecture. In this work, we define ring homomorphisms on $F[x,y,z]$, given in (2), extending the Nagata automorphism which are wild automorphisms under some assumptions, see Corollary 1). We also relate the condition of non-vanishing of their Jacobian determinant with a certain linear partial differential equation, see Theorem 1. As a consequence, we prove that those ring homomorphisms on $F[x,y,z]$ with nonzero Jacobian determinant, are in fact automorphisms, see Theorem 2.
2. Nagata polynomial map
For a polynomial $\varphi\in F[x,y,z]$, we set
(1) $\begin{split}f(x,y,z)&=x-2y\varphi-z\varphi^{2}\,,\\ g(x,y,z)&=y+z\varphi\,,\\ h(x,y,z)&=z\,.\end{split}$
The triplet $(f,g,h)$ induces a ring homomorphism on $F[x,y,z]$. Note that $(f,g,h)$ is the identity when one takes $\varphi=0$, so $(f,g,h)$ may be seen as a perturbation of the identity. If $\varphi=xz+y^{2}$ then $(f,g,h)$ is the Nagata automorphism. The Milnor number of $(f,g,h)$ is the dimension
$\mu(f,g,h)=\dim_{F}\frac{F[x,y,z]}{\langle f,g,h\rangle}$
as vector space over $F$.
Lemma 1.
If $(f,g,h)$ is a ring homomorphism as in (2), then $\mu(f,g,h)=1$.
Proof.
One has,
$y$ $=(y+z\varphi)-\varphi z=g-\varphi z,$
$x$ $=(x-2y\varphi-z\varphi^{2})+2(y+z\varphi)\varphi-\varphi^{2}z=f+2\varphi g-\varphi^{2}z,$
hence we get an equality $\langle f,g,h\rangle=\langle x,y,z\rangle$ of polynomial ideals, and this implies that $\mu(f,g,h)=1$. ∎
Remark 1.
Note that $(f,g,h)$ is not always an automorphism for any $\varphi$ despite the fact that $\mu(f,g,h)=1$ by Lemma 1. For instance, if $\varphi=x$ or $\varphi=y$, then $(f,g,h)$ is not an automorphism. In the first case, $(f,g,h)$ sends both points $(0,0,1)$ and $(-1,1,1)$ to $(0,0,1)$. Theorem 2 provides a necessary and sufficient condition on $\varphi$ such that $(f,g,h)$ is an automorphism.
Proposition 1.
The Jacobian matrix $J(f,g,h)$ has nonzero determinant in $F$ if and only if $\varphi$ satisfies the linear PDE:
$-2y\varphi_{x}+z\varphi_{y}=0\,.$
Proof.
We have
$f_{x}$ $=1-2y\varphi_{x}-2z\varphi\varphi_{x},$
$f_{y}$ $=-2\varphi_{x}-2y\varphi_{y}-2z\varphi\varphi_{y},$
$g_{x}$ $=z\varphi_{x},$
$g_{y}$ $=1+z\varphi_{y}\,.$
Hence the Jacobian matrix $J(f,g,h)$ is
$\begin{pmatrix}1-2y\varphi_{x}-2z\varphi\varphi_{x}&-2\varphi_{x}-2y\varphi_{y}-2z\varphi\varphi_{y}&f_{z}\\ z\varphi_{x}&1+z\varphi_{y}&g_{z}\\ 0&0&1\end{pmatrix}$
whose determinant is
$\begin{vmatrix}1-2y\varphi_{x}-2z\varphi\varphi_{x}&-2\varphi_{x}-2y\varphi_{y}-2z\varphi\varphi_{y}\\ z\varphi_{x}&1+z\varphi_{y}\end{vmatrix}=1-2y\varphi_{x}+z\varphi_{y}\,.$
If $-2y\varphi_{x}+z\varphi_{y}=0$ then the determinant of $J(f,g,h)$ is $1$. Conversely, if the determinant of $J(f,g,h)$ is a constant $c\in F-\{0\}$, then
$1-2y\varphi_{x}+z\varphi_{y}=c\,.$
However this equality is only possible when $c=1$, therefore $-2y\varphi_{x}+z\varphi_{y}=0$. ∎
Proposition 2.
Let $p\in F[t_{1},t_{2}]$ and set $\varphi=p(xz+y^{2},z)$ in $F[x,y,z]$. Then the ring homomorphism $(f,g,h)$ given in (2) is an automorphism on $F[x,y,z]$ with inverse:
(2) $\begin{split}f^{\prime}(x,y,z)&=x+2y\varphi-z\varphi^{2},\\ g^{\prime}(x,y,z)&=y-z\varphi,\\ h^{\prime}(x,y,z)&=z\,.\end{split}$
Proof.
We have
$\varphi(f^{\prime},g^{\prime},h^{\prime})$ $=p(f^{\prime}z+g^{\prime 2},z)$
$=p\Big{(}(x+2y\varphi-z\varphi^{2})z+(y-z\varphi)^{2},z\Big{)}$
$=p(xz+y^{2},z)$
$=\varphi\,,$
thus $\varphi(f^{\prime},g^{\prime},h^{\prime})=\varphi$. Hence,
$f(f^{\prime},g^{\prime},h^{\prime})$ $=f^{\prime}-2g^{\prime}\varphi(f^{\prime},g^{\prime},h^{\prime})-z\varphi^{2}(f^{\prime},g^{\prime},h^{\prime})$
$=(x+2y\varphi-z\varphi^{2})-2(y-z\varphi)\varphi-z\varphi^{2}$
$=x\,.$
On the other hand, we have
$g(f^{\prime},g^{\prime},h^{\prime})$ $=g^{\prime}+z\varphi(f^{\prime},g^{\prime},h^{\prime})$
$=(y-z\varphi)+z\varphi$
$=y\,.$
Therefore, the composite $(f,g,h)\circ(f^{\prime},g^{\prime},h^{\prime})$ is the identity. Similarly, one verifies that $(f^{\prime},g^{\prime},h^{\prime})\circ(f,g,h)$ is also the identity. ∎
For a polynomial $\psi$, we denote by $\overline{\psi}$ the highest degree homogeneous component of $\psi$.
Proposition 3.
Let $\varphi\in F[x,y,z]$ and suppose that $(f,g,h)$ is an automorphism on $F[x,y,z]$ of the form (2). If one the partial derivatives $\overline{\varphi}_{x},\overline{\varphi}_{y}$ is nonzero, then the automorphism $(f,g,h)$ is wild.
Proof.
Let $\overline{f},\overline{g},\overline{h}$ be the highest homogeneous component of $f,g,h$, respectively. Since $\overline{\varphi}_{x}$ or $\overline{\varphi}_{y}$ is nonzero, we have $\deg(\overline{\varphi})>0$, so
$\begin{split}\overline{f}&=-z\overline{\varphi}^{2}\,,\\ \overline{g}&=z\overline{\varphi}\,,\\ \overline{h}&=z\,.\end{split}$
Hence,
$\overline{f}_{x}\overline{g}_{y}-\overline{f}_{y}\overline{g}_{x}$ $=0$
$\overline{f}_{y}\overline{g}_{z}-\overline{f}_{z}\overline{g}_{y}$ $=-z\,\overline{\varphi}^{2}\overline{\varphi}_{y}$
$\overline{f}_{x}\overline{g}_{z}-\overline{f}_{z}\overline{g}_{x}$ $=-z\,\overline{\varphi}^{2}\overline{\varphi}_{x}$
$\overline{g}_{x}\overline{h}_{y}-\overline{g}_{y}\overline{h}_{x}$ $=0$
$\overline{g}_{y}\overline{h}_{z}-\overline{g}_{z}\overline{h}_{y}$ $=z\,\overline{\varphi}_{y}$
$\overline{g}_{x}\overline{h}_{z}-\overline{g}_{z}\overline{h}_{x}$ $=z\,\overline{\varphi}_{x}$
$\overline{f}_{x}\overline{h}_{y}-\overline{f}_{y}\overline{h}_{x}$ $=0$
$\overline{f}_{y}\overline{h}_{z}-\overline{f}_{z}\overline{h}_{y}$ $=-2z\,\overline{\varphi}_{y}$
$\overline{f}_{x}\overline{h}_{z}-\overline{f}_{z}\overline{h}_{x}$ $=-2z\,\overline{\varphi}_{x}\,.$
By [3, Corollary 2] or [2, Lemma 1], we deduce that $\overline{f},\overline{g},\overline{h}$ are mutually algebraically independent. On the other hand, from the hypothesis $\overline{\varphi}\in F[x,y,z]$ is a homogeneous polynomial of positive degree such that $\overline{\varphi}\not\in F[z]$, it is not difficult to see that none of the polynomials $\overline{f},\overline{g},\overline{h}$ belongs to the subalgebra generated by the other two polynomials. Therefore, by [2, Corollary 8] $(f,g,h)$ is wild. ∎
We give the following definition in order to determine the highest weighted homogeneous component of the polynomial $\varphi=p(xz+y^{2},z)$.
Definition 1.
Let us fix algebraically independent indeterminates $t_{1},\dots,t_{n}$ over $F$. For $k\in\mathbb{Z}^{n}_{\geq 1}$, we write $t^{k}$ to denote the monomial $t_{1}^{k_{1}}\cdots t_{n}^{k_{n}}$. A polynomial $h\in F[t_{1},\dots,t_{n}]$ can be written as $h=\sum_{k}a_{k}t^{k}$.
  1. (1)
    The support of $h$, denoted by $\mathrm{supp}(h)$, is defined as the set
    $\mathrm{supp}(h)=\{k\in\mathbb{Z}^{n}_{\geq 0}:a_{k}\neq 0\}\,.$
  2. (2)
    Let $v=(v_{1},\dots,v_{n})\in\mathbb{Z}^{n}_{\geq 1}$. The degree of $h$ with respect to $v$, denoted by $d_{v}(h)$, is the maximum of the scalar product $\langle v,k\rangle$ such that $a_{k}\neq 0$. We denote by $h^{v}$ the sum of those terms $a_{k}t^{k}$ such that $\langle v,k\rangle=d_{v}(p)$. In particular, if $v=(1,\dots,1)$ then $\deg(h)=d_{v}(h)$ and $h^{v}=\overline{h}$.
  3. (3)
    We say that $h$ is weighted homogeneous with respect to $v$ when $h^{v}=h$.
Lemma 2.
Let $p\in F[t_{1},t_{2}]$ and $v=(2,1)$. If $\varphi=p(xz+y^{2},z)\in F[x,y,z]$, then $\deg(\varphi)=d_{v}(p)$ and $\overline{\varphi}=p^{v}(xz+y^{2},z)$. In particular, if $p$ is weighted homogeneous with respect to $v=(2,1)$, then $\varphi$ is homogeneous.
Proof.
Let us write the polymomial $p$ as $p=\sum_{k}a_{k}t^{k}$. One has
$\varphi=\sum_{k}a_{k}(xz+y^{2})^{k_{1}}z^{k_{2}}=\sum_{k}a_{k}\sum_{n=0}^{k_{1}}\binom{k_{1}}{n}x^{k_{1}-n}y^{2n}z^{k_{1}+k_{2}-n}\,.$
From the last equality, we observe that
(3) $\mathrm{supp}(\varphi)=\Big{\{}(k_{1}-n,2n,k_{1}+k_{2}-n):k\in\mathrm{supp}(p)\,\textnormal{and}\,\,0\leq n\leq k_{1}\Big{\}}\,.$
Therefore,
$\deg(\varphi)$ $=\max\Big{\{}k_{1}-n+2n+k_{1}+k_{2}-n:(k_{1},k_{2})\in\mathrm{supp}(p)\,\textnormal{and}\,0\leq n\leq k_{1}\Big{\}}$
$=\max\Big{\{}2k_{1}+k_{2}:(k_{1},k_{2})\in\mathrm{supp}(p)\Big{\}}$
(4) $=d_{v}(p)\,.$
Hence, by the equalities given in (3) and (4), the highest degree homogeneous component of $\varphi$ is computed as follows
$\overline{\varphi}=$ $\sum_{\begin{subarray}{c}k\in\mathrm{supp}(p)\\ 0\leq n\leq k_{1}\\ 2k_{1}+k_{2}=d_{v}(p)\end{subarray}}a_{k}\sum_{n=0}^{k_{1}}\binom{k_{1}}{n}x^{k_{1}-n}y^{2n}z^{k_{1}+k_{2}-n}$
$=$ $\sum_{\begin{subarray}{c}k\in\mathrm{supp}(p)\\ 0\leq n\leq k_{1}\\ 2k_{1}+k_{2}=d_{v}(p)\end{subarray}}a_{k}\left(\sum_{n=0}^{k_{1}}\binom{k_{1}}{n}x^{k_{1}-n}z^{k_{1}-n}y^{2n}\right)z^{k_{2}}$
$=$ $\sum_{\begin{subarray}{c}k\in\mathrm{supp}(p)\\ 0\leq n\leq k_{1}\\ 2k_{1}+k_{2}=d_{v}(p)\end{subarray}}a_{k}\left(\sum_{n=0}^{k_{1}}\binom{k_{1}}{n}(xz)^{k_{1}-n}y^{2n}\right)z^{k_{2}}$
$=$ $\sum_{\begin{subarray}{c}k\in\mathrm{supp}(p)\\ 2k_{1}+k_{2}=d_{v}(p)\end{subarray}}\left(xz+y^{2}\right)^{k_{1}}z^{k_{2}}$
$=$ $p^{v}(xz+y^{2},z)\,,$
and it proves the assertion. ∎
Example 1.
If $\varphi(x,y,z)=p(xz+y^{2},z)$, then it is not always true that $\overline{\varphi}(x,y,z)=\overline{p}(xz+y^{2},z)$. Indeed, consider $p(t_{1},t_{2})=t_{1}^{2}-t_{2}^{3}+t_{1}t_{2}^{2}\in F[s,t]$, then by a immediate computation one has
$\overline{\varphi}=\left(xz+y^{2}\right)^{2}+\left(xz+y^{2}\right)z^{2}\neq-z^{3}+\left(xz+y^{2}\right)z^{2}=\overline{p}(xz+y^{2},z)\,.$
However, it is easy to verify that if $v=(2,1)$, then $d_{v}(p)=4$ and $p^{v}=t_{1}^{2}+t_{1}t_{2}^{2}$. Moreover, by direct computation one has
$\overline{\varphi}=\left(xz+y^{2}\right)^{2}+\left(xz+y^{2}\right)z^{2}=p^{v}(xz+y^{2},z)\,.$
Corollary 1.
Let $p\in F[t_{1},t_{2}]$ and set $\varphi=p(xz+y^{2},z)$ in $F[x,y,z]$. If the partial derivative $(p^{v})_{t_{1}}$ is nonzero, then the corresponding ring homomorphism $(f,g,h)$ given in (2) is a wild automorphism.
Proof.
By Proposition 2, $(f,g,h)$ is an automorphism and Lemma 2 establishes the equality $\overline{\varphi}=p^{v}(xz+y^{2},z)$. Hence the fact that $(p^{v})_{t_{1}}\neq 0$ implies that $\overline{\varphi}_{x}\neq 0$ and $\overline{\varphi}_{y}\neq 0$. By Proposition 3, the automorphism $(f,g,h)$ is wild. ∎
Example 2.
Let $\varphi=(xz+y^{2})^{n}$ with $n\geq 1$. Corollary 1 implies that
$f(x,y,z)$ $=x-2y(xz+y^{2})^{n}-z(xz+y^{2})^{2n},$
$g(x,y,z)$ $=y+z(xz+y^{2})^{n},$
$h(x,y,z)$ $=z,$
is a wild automorphism.
Example 3.
Let $\varphi(x,y,z)=p(xz+y^{2},z)$, where $p(t_{1},t_{2})=t_{1}^{2}-t_{2}^{3}+t_{1}t_{2}^{2}\in F[s,t]$ as in Example 1. Since $p^{v}=t_{1}^{2}+t_{1}t_{2}^{2}$ and $(p^{v})_{t_{1}}=2t_{1}+t_{1}^{2}\neq 0$, Corollary 1 implies that corresponding ring homomorphism $(f,g,h)$ is a wild automorphism.
3. Statement of the main results
The main results in this work are the following:
Theorem 1.
The linear PDE
(5) $-2y\varphi_{x}+z\varphi_{y}=0$
has as general solution in $F[x,y,z]$, polynomials that can be written as:
$\varphi=p(xz+y^{2},z)\,,$
where $p\in F[t_{1},t_{2}]$.
Theorem 2.
Let $\varphi\in F[x,y,z]$ and set
$f(x,y,z)$ $=x-2y\varphi-z\varphi^{2}\,,$
$g(x,y,z)$ $=y+z\varphi\,,$
$h(x,y,z)$ $=z\,.$
Then, the determinant of the Jacobian matrix $J(f,g,h)$ is a nonzero constant in $F$ if and only if $(f,g,h)$ is an automorphism on $F[x,y,z]$. In this case, there exists a polynomial $p\in F[t_{1},t_{2}]$ such that $\varphi=p(xz+y^{2},z)$, moreover the inverse of $(f,g,h)$ is given in (2).
The proofs of these results are given after the following previous results.
Lemma 3.
Let $\varphi=\varphi^{0}+\varphi^{1}+\cdots+\varphi^{n}$ be a solution of the linear PDE:
(6) $A(x,y,z)\varphi_{x}+B(x,y,z)\varphi_{y}+C(x,y,x)\varphi_{z}=0\,,$
where each $\varphi^{k}\in F[x,y.z]$ is the homogeneous component of $\varphi$ of degree $k$ for $0\leq k\leq n$, and $A,B,C$ are homogeneous polynomials of the same degree. Then each homogeneous component $\varphi^{k}$ is also a solution of the equation (6).
Proof.
Replacing $\varphi=\sum_{k=0}^{n}\varphi^{k}$ into the linear PDE (5), we obtain:
$\sum_{k=0}^{n}\Big{(}A(x,y,z)\varphi^{k}_{x}+B(x,y,z)\varphi^{k}_{y}+C(x,y,x)\varphi^{k}_{z}\Big{)}=0$
which is a sum of homogeneous polynomials of different degrees, therefore:
$A(x,y,z)\varphi^{k}_{x}+B(x,y,z)\varphi^{k}_{y}+C(x,y,x)\varphi^{k}_{z}=0\,\,\,\textnormal{for all}\,\,k=0,1,\dots,n\,.$
By the previous lemma, to find the general solution of the equation (5), it is enough to deal with its homogeneous polynomials components. Let $\varphi\in F[x,y,z]$ be a homogeneous polynomial of degree $n$, if we consider the canonical isomorphism of rings $F[x,y,z]\cong F[x,y][z]$, then we can suppose that $\varphi$ is written as:
(7) $\varphi=\varphi^{n}+\varphi^{n-1}z+\varphi^{n-2}z^{2}+\dots+\varphi^{1}z^{n-1}+\varphi^{0}z^{n}=\sum_{k=0}^{n-1}\varphi^{n-k}z^{k}+\varphi^{0}z^{n}$
with $\varphi^{k}\in F[x,y]$ homogeneous polynomials of degree $k$ for $k=0,1,\dots,n$.
Lemma 4.
Let $\varphi$ be a homogeneous polynomial in $F[x,y,z]$ of degree $n$ that is a solution of the linear PDE (5). Then, $\varphi^{n}=ay^{n}$ and $\varphi^{1}=bx$, where $a$ and $b$ are constants in $F$, and moreover
$2y\varphi^{n-k-1}_{x}=\varphi^{n-k}_{y}$
for all $k=0,1,\dots,n-2$.
Proof.
Let us derivate $\varphi$ with respect to $x$ and $y$:
$\varphi_{x}$ $=\sum_{k=0}^{n-1}\varphi^{n-k}_{x}z^{k}=\varphi_{x}^{n}+\sum_{k=1}^{n-1}\varphi^{n-k}_{x}z^{k}\,,$
$\varphi_{y}$ $=\sum_{k=0}^{n-1}\varphi^{n-k}_{y}z^{k}=\sum_{k=0}^{n-2}\varphi^{n-k}_{y}z^{k}+\varphi^{1}_{y}z^{n-1}\,.$
Substituting these equalities in the equation (5), we have
$-2y\varphi_{x}^{n}+\sum_{k=1}^{n-1}-2y\varphi^{n-k}_{x}z^{k}+\sum_{k=0}^{n-2}\varphi^{n-k}_{y}z^{k+1}+\varphi^{1}_{y}z^{n}=0\,,$
and by index rescaling, we get
$-2y\varphi_{x}^{n}+\sum_{k=0}^{n-2}-2y\varphi^{n-k-1}_{x}z^{k+1}+\sum_{k=0}^{n-2}\varphi^{n-k}_{y}z^{k+1}+\varphi^{1}_{y}z^{n}=0\,.$
Hence,
$-2y\varphi_{x}^{n}+\sum_{k=0}^{n-2}\left(-2y\varphi^{n-k-1}_{x}+\varphi^{n-k}_{y}\right)z^{k+1}+\varphi^{1}_{y}z^{n}=0\,.$
This equation gives the equalities $\varphi^{n}_{x}=0$ and $\varphi^{1}_{y}=0$, so $\varphi^{n}=ay^{n}$ and $\varphi^{1}=bx$ for some constants $a,b\in F$. Moreover,
(8) $2y\varphi^{n-k-1}_{x}=\varphi^{n-k}_{y}\,\,\,\textnormal{para todo}\,\,\,k=0,1,\dots,n-2\,.$
The following result, will enable us to know how a solution can be written of the equation (5).
Proposition 4.
Let $\varphi\in F[x,y,z]$ be a homogeneous solution of the linear PDE (5). We have the following assertions:
  1. ($a$)
    If $\varphi$ has degree $2n$, then
    $\varphi=a(xz+y^{2})^{n}+z\psi$
    where $a\in F$ and $\psi\in F[x,y,z]$ is a homogeneous polynomial which is also a solution of the equation (5) of degree $12n-$ 1.
  2. ($b$)
    If $\varphi$ has degree $12n+$ 1, then
    $\varphi=z\psi$
    where $\psi\in F[x,y,z]$ is homogeneous of degree $2n$ which is also a solution of (5) of degree $2n$.
Proof.
($a$). Let us suppose that:
(9) $\varphi=\varphi^{2n}+\varphi^{2n-1}z+\cdots+\varphi^{1}z^{2n-1}+\varphi^{0}z^{2n}$
with $\varphi^{2n-k}$ be a homogeneous polynomial in $F[x,y]$ of degree $2n-k$ for $0\leq k\leq 2n$. By Lemma 4, we get the following equalities:
(10) $\varphi^{2n}$ $=ay^{2n}\,\,\,\,\textnormal{and}\,\,\,\,\varphi^{1}=bx$
(11) $2y\varphi^{2n-k-1}_{x}$ $=\varphi^{2n-k}_{y}\,\,\,\textnormal{para todo}\,\,\,k=0,1,\dots,2n-2\,.$
We claim that there are constants $a,a^{2n-k}_{j}$, with $1\leq k\leq n$ y $10\leq j\leq k-$ 1, such that:
(12) $\varphi^{2n-k}$ $=a\binom{n}{k}y^{2n-2k}x^{k}+\sum_{j=0}^{k-1}a^{2n-k}_{k-1-j}y^{2n-2k+1+j}x^{k-1-j}$
Indeed, taking $k=0$, from these equalities (10) and (11) we find that:
$2y\varphi^{2n-1}_{x}=\varphi^{2n}_{y}=2nay^{2n-1}$
so $\varphi_{x}^{2n-1}=nay^{2n-2}$. Hence, by integrating with respect to $x$, we obtain
$\varphi^{2n-1}=nay^{2n-2}x+a^{2n-1}_{0}y^{2n-1}\,,$
which can be written as
$\varphi^{2n-1}=a\binom{n}{1}y^{2n-2}x+a^{2n-1}_{0}y^{2n-1}\,,$
thus (12) is true for $k=1$. Suppose we have (12). Taking the derivative with respect to $y$, we have
$\varphi^{2n-k}_{y}=2a\binom{n}{k}(n-k)y^{2n-2k-1}x^{k}+\sum_{j=0}^{k-1}\,.(2n-2k+1+j)a^{2n-k}_{k-1-j}y^{2n-2k+j}x^{k-1-j}$
From the equality (11), we get
$\varphi^{2n-k-1}_{x}=a\binom{n}{k}(n-k)y^{2n-2k-2}x^{k}+\sum_{j=0}^{k-1}\frac{(2n-2k+1+j)a^{2n-k}_{k-1-j}}{2}y^{2n-2k-1+j}x^{k-1-j}\,.$
Integrating with respect to $x$, there is a constant $a^{2n-k-1}_{0}\in F$ such that
$\varphi^{2n-k-1}=a\binom{n}{k+1}y^{2n-2k-2}x^{k+1}+\sum_{j=0}^{k-1}\frac{(2n-2k+1+j)a^{2n-k}_{k-1-j}}{2(k-1-j)}y^{2n-2k-1+j}x^{k-j}+a^{2n-k-1}_{0}y^{2n-k-1}\,.$
Defining $a^{2n-k-1}_{k-j}=\frac{(2n-2k+1+j)a^{2n-k}_{k-1-j}}{2(k-1-j)}$, we get
$\varphi^{2n-k-1}=a\binom{n}{k+1}y^{2n-2k-2}x^{k+1}+\sum_{j=0}^{k}a^{2n-k-1}_{k-j}y^{2n-2k-1+j}x^{k-j}$
as wished, and this completes the recurrence. Now, replacing the expressions obtained in (10) and (12) in the equation (9), we have
$\varphi$ $=\sum_{k=0}^{n}\varphi^{2n-k}z^{k}+\sum_{k=n+1}^{2n}\varphi^{2n-k}z^{k}$
$=\varphi^{2n}+\sum_{k=1}^{n}\varphi^{2n-k}z^{k}+\sum_{k=n+1}^{2n}\varphi^{2n-k}z^{k}$
$=ay^{2n}+\sum_{k=1}^{n}\left(a\binom{n}{k}y^{2n-2k}x^{k}+\sum_{j=0}^{k-1}a^{2n-k}_{k-1-j}y^{2n-2k+1+j}x^{k-1-j}\right)z^{k}+\sum_{k=n+1}^{2n}\varphi^{2n-k}z^{k}$
$=a\sum_{k=0}^{n}\binom{n}{k}y^{2n-2k}x^{k}z^{k}+\sum_{k=1}^{n}\left(\sum_{j=0}^{k-1}a^{2n-k}_{k-1-j}y^{2n-2k+1+j}x^{k-1-j}\right)z^{k}+\sum_{k=n+1}^{2n}\varphi^{2n-k}z^{k}\,.$
From the last equality, we deduce that
$\varphi=a(xz+y^{2})^{n}+z\psi,$
where $\psi\in F[x,y,z]$ is a homogeneous polynomial of degree equal to $12n-$ 1. By the linearity of the solutions of the equation (5), we obtain that $z\psi$ is a solution, hence $\psi$ is also a solution. ($b$) Let us suppose that:
(13) $\varphi=\varphi^{2n+1}+\varphi^{2n}+\cdots+\varphi^{1}z^{2n}+\varphi^{0}z^{2n+1}\,,$
where $\varphi^{2n-k}$ is a homogeneous polynomial in $F[x,y]$ of degree equal to $2n+1-k$, for all $10\leq k\leq 2n+$ 1. By Lemma 4, we get the following equalities:
(14) $\varphi^{2n+1}$ $=ay^{2n+1}$
(15) $2y\varphi^{2n-k}_{x}$ $=\varphi^{2n+1-k}_{y}\,\,\,\textnormal{para todo}\,\,\,k=0,1,\dots,2n-1\,.$
As in the even case, we claim that there are constants $a,a_{j}^{2n-k}$, with $10\leq k\leq n-$ 1 and $0\leq j\leq k$, such that:
(16) $\varphi^{2n-k}=\frac{(2n-2k+1)\cdots(2n-1)(2n+1)}{2^{k}\cdot(k+1)!}ay^{2n-2k-1}x^{k+1}+\sum_{j=0}^{k}b^{2n-k}_{k-j}y^{2n-2k+j}x^{k-j}\,.$
Replacing $k=0$ in the equation (14) and joining with (15), we have
$2y\varphi^{2n}_{x}=\varphi^{2n+1}_{y}=(2n+1)ay^{2n}\,,$
hence $\varphi_{x}^{2n}=(2n+1)ay^{2n-1}$. Integrating with respect to $x$, we obtain
$\varphi^{2n}=(2n+1)ay^{2n-1}x+b^{2n}_{0}y^{2n}$
thus (16) is true for $k=1$. Suppose we have (16). Taking the derivative with respect to $y$, we have
$\varphi^{2n-k}_{y}=\frac{(2n-2k-1)\cdots(2n-1)(2n+1)}{2^{k}\cdot(k+1)!}ay^{2n-2k-2}x^{k+1}+\sum_{j=0}^{k}(2n-2k+j)b^{2n-k}_{k-j}y^{2n-2k+j-1}x^{k-j}\,.$
From the equality (15), we get
$\varphi^{2n-k-1}_{x}=\frac{(2n-2k-1)\cdots(2n-1)(2n+1)}{2^{k+1}\cdot(k+1)!}ay^{2n-2k-3}x^{k+1}+\sum_{j=0}^{k}\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2}y^{2n-2k+j-2}x^{k-j}\,.$
Integrating with respect to $x$, there is a constant $b^{2n-k-1}_{0}\in F$ such that $\varphi^{2n-k-1}$ is equal to
$\frac{(2n-2k-1)\cdots(2n-1)(2n+1)}{2^{k+1}\cdot(k+2)!}ay^{2n-2k-3}x^{k+2}+\sum_{j=0}^{k}\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2(k-j+1)}y^{2n-2k+j-2}x^{k-j+1}+b^{2n-k-1}_{0}y^{2n-k-1}\,.$
Defining $b^{2n-k-1}_{k-j}=\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2(k-j+1)}$, we get
$\varphi^{2n-k-1}=\frac{(2n-2k+-1)\cdots(2n-1)(2n+1)}{2^{k+1}\cdot(k+2)!}ay^{2n-2k-3}x^{k+2}+\sum_{j=0}^{k+1}\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2(k-j+1)}y^{2n-2k+j-2}x^{k-j+1}$
as wished, and this completes the recurrence. In particular, replacing $k=n-1$ in (16), we obtain:
$\varphi^{n+1}=\frac{1\cdot 3\cdots(2n+1)}{2^{n-1}\cdot n!}ayx^{n}+\sum_{j=0}^{n-1}b^{n+1}_{n-1-j}y^{j+2}x^{n-1-j}\,.$
We solve the equation (15) for $k=n$ and we get:
$-2y\varphi^{n}_{x}$ $=\varphi^{n+1}_{y}\,,$
(17) $-2y\varphi^{n}_{x}$ $=\frac{1\cdot 3\cdots(2n+1)}{2^{n-1}\cdot n!}ax^{n}+\sum_{j=0}^{n-1}(j+2)b^{n+1}_{n-1-j}y^{j+1}x^{n-1-j}\,.$
From the equality (17), we deduce that $a=0$ and as consequence from this we obtain that $\varphi^{2n+1}=0$. Therefore,
$\varphi=z\varphi^{2n}+z^{2}\varphi^{2n-1}+\cdots+z^{2n}\varphi^{1}+\varphi^{0}z^{2n+1}=z\psi\,,$
where $\psi\in F[x,y,z]$ is a homogeneous polynomial of degree $2n$, and as in the previous case, we obtain that $\psi$ is also a solution. ∎
4. Proof of the main results
Proof of Theorem 1.
Suppose that $\varphi=p(xz+y^{2},z)$, where $p$ is a polynomial in $F[t_{1},t_{2}]$. Replacing the partial derivatives
$\varphi_{x}$ $=z\cdot p_{s}(xz+y^{2},z)\,,$
$\varphi_{y}$ $=2y\cdot p_{s}(xz+y^{2},z)\,.$
in the linear PDE (5), one verifies that $\varphi=p(xz+y^{2},z)$ is a solution. Reciprocally, suppose that $\varphi\in F[x,y,z]$ is a solution of the equation (5). We shall prove that $\varphi$ has the form $\varphi=p(xz+y^{2},z)$, where $p$ is a polynomial in $F[t_{1},t_{2}]$. By Lemma 3, we can assume that $\varphi\in F[x,y,z]$ is a homogeneous polynomial. More precisely, we shall prove by recurrence the following:
(18) $\varphi=\left\{\begin{array}[]{rcl}\sum_{k=0}^{n}a_{2k}z^{2k}(xz+y^{2})^{2n-2k}&\mbox{if}&\deg\varphi=2n\,,\\ &&\\ \sum_{k=1}^{n}a_{2k-1}z^{2k-1}(xz+y^{2})^{2n-2k}&\mbox{if}&\deg\varphi=2n-1\,.\end{array}\right.$
Indeed, applying Proposition 4 for $n=1$, we obtain:
$\varphi=\left\{\begin{array}[]{ccl}a(xz+y^{2})+z\psi&\mbox{if}&\deg\varphi=2\,,\\ &&\\ z\psi_{1}&\mbox{if}&\deg\varphi=1\,,\end{array}\right.$
where $\psi$ and $\psi_{1}$ are homogeneous polynomials with $\deg\psi=1$ and $\deg\psi_{1}=0$. Then there are constants $b,c\in F$ such that $\psi=cz$ and $\psi_{1}=b\in F$, so (18) is true for $n=1$.
Let us suppose that (18) holds for homogeneous polynomial solutions of degree $2n$ or $12n-$ 1. If $\deg\varphi=2n+2$, then by Proposition 4 ($a$), we obtain that $\varphi=a(xz+y^{2})^{2n+2}+z\psi$, where $\psi$ is a solution of the equation (5). By the inductive hypothesis, we have
$\psi=\sum_{k=1}^{n}a_{2k+1}z^{2k-1}(xz+y^{2})^{2n-2k}\,.$
Hence,
$\varphi$ $=a(xz+y^{2})^{2n+2}+z\psi$
$=a(xz+y^{2})^{2n+2}+z\left(\sum_{k=1}^{n}a_{2k+1}z^{2k-1}(xz+y^{2})^{2n-2k}\right)\,.$
If $\deg\varphi=2n+1$, then by Proposition 4 ($b$), we obtain that $\varphi=z\psi$, where $\psi$ is a solution of the equation (5) with $\deg\psi=2n$. By the inductive hypothesis, one has
$\psi=\sum_{k=0}^{n}a_{2k}z^{2k}(xz+y^{2})^{2n-2k}\,.$
Hence,
$\varphi$ $=z\psi$
$=z\left(\sum_{k=0}^{n}a_{2k}z^{2k}(xz+y^{2})^{2n-2k}\right)$
$=\sum_{k=0}^{n}a_{2k}z^{2k+1}(xz+y^{2})^{2n-2k}$
$=\sum_{k=1}^{n+1}a_{2k-2}z^{2k-1}(xz+y^{2})^{2(n+1)-2k}\,,$
which completes the proof. ∎
Proof of the Theorem 2.
One implication is straightforward. Suppose that $\det J(f,g,h)$ is a nonzero constant. By Proposition 1, $\varphi$ satisfies the linear PDE (5). By Theorem 1, the solution of the equation (5) has the form $\varphi=p(xz+y^{2},z)$ where $p$ is a polynomial in $F[t_{1},t_{2}]$. Finally, by Proposition 2, $(f,g,h)$ is an automorphism and its inverse is given in (2). ∎
5. Conclusions
Theorem 2 says that the automorphisms of the form (2) satisfy the well know Jacobian conjecture via the approach of solving first order partial differential equations. A further problem is whether those automorphisms might be naturally extended on polynomial rings with more than three variables satisfying the aforementioned result.